# NextStep-1: Toward Autoregressive Image Generation with Continuous Tokens at Scale

### v1

@@ 1 Introduction
However, most existing autoregressive approaches for text-to-image generation (Sun et al., 2024b; Wang et al., 2024b; Sun et al., 2024a; Chen et al., 2025b; Tong et al., 2024; Dong et al., 2024) either rely on heavy diffusion models or adopt vector quantization (VQ) (Zheng et al., 2022; Eslami et al., 2021; Yu et al., 2023) to tokenize images into discrete visual tokens, which encounter limitations including exposure bias (Han et al., 2025) and suboptimal image tokenization (Li et al., 2024c). While recent efforts with continuous latent representations (Li et al., 2024c; Fan et al., 2024) have shown promise, a significant performance gap persists between autoregressive models and state-of-the-art diffusion methods (Podell et al., 2024; Esser et al., 2024; Labs, 2024), particularly in image quality and consistency.
@@ 2.2 Model Architecture
The latents from the image tokenizer are pixel-shuffled into a more compact sequence. This is achieved by applying a space-to-depth transformation with a $\times2$ kernel, which flattens $\times2$ spatial latents into the channel dimension. For example, this converts the latents of a $\times256$ image into $\times16$ grid of 64-channel tokens. This grid is then flattened into a 1D sequence of 256 tokens to serve as input for the following Causal Transformer.
@@ 3.2 Image-Text Pair Data
    Data Sourcing: We collected a large-scale dataset from diverse sources, including web data, multi-task VQA data and text-rich documents.
@@ 3.3 Instruction-Guided Image-to-Image Data
For visual perception and controllable image generation tasks, we synthesized 1M samples by applying the annotator of ControlNet (Zhang et al., 2023b) to a part of our high-quality image-text pair data. For image restoration and general image editing, we collected 3.5M samples, comprising data from GPT-Image-Edit (Wang et al., 2025c), Step1X-Edit (Liu et al., 2025), and a proprietary in-house dataset. Following Step1X-Edit (Liu et al., 2025), all editing data were subjected to a rigorous VLM-based filtering pipeline that assessed both image-pair quality, rationality, consistency, and instruction alignment, resulting in about 1M high-quality instruction-guided image-to-image data for training.
@@ 4.1 Training Image Tokenizer
Our image tokenizer is initialized from the Flux.1-dev VAE (Labs, 2024), selected for its strong reconstruction performance. We fine-tune this model on the image-text dataset detailed in Section˜3.2 to adapt it to our specific data distribution. For optimization, we employ the AdamW optimizer (Loshchilov and Hutter, 2019) with $(\beta_{1}=0.9,\beta_{2}=0.95,\varepsilon=1\times 10^{-8})$ for its convergence stability. The model is trained for 50K steps with a total batch size of 512, using a constant learning rate of $1\times 10^{-5}$ preceded by linear warm-up of 1000 steps.
@@ 4.3 Post-Training
PixArt-Alpha (Chen et al., 2024) 0.48 - - 71.11
@@ 5.1 Performance of Text-to-Image Generation
PixArt-Alpha (Chen et al., 2024) 0.45 0.50 0.48 0.49 0.56 0.34 0.47 0.63
@@
GPT-Image-Edit (Wang et al., 2025c) - - 7.24 - - - 3.80
@@ 6.1 What Governs Image Generation: the AR Transformer or the FM Head?
A key architectural distinction of our framework is its direct, autoregressive modeling of continuous image tokens using a flow matching objective. Prevailing autoregressive models for image generation  (Sun et al., 2023, 2024b; Dong et al., 2024; Zhou et al., 2025; Chen et al., 2025a) typically rely on heavy, diffusion models for a entire image: an autoregressive model first produces a semantic embedding, which is then used to condition a diffusion model that generates an entire image in a single denoising process. In contrast, our model autoregressively generates the image patch-by-patch, modeling the distribution of each patch with a significantly more lightweight flow matching model. We argue that this establishes our framework under the pure autoregressive paradigm with next-token prediction (NTP) modeling, rather than a diffusion model merely orchestrated by a Transformer.
@@ 6.2 Tokenizer is the Key to Image Generation
where $v_{\theta}(x|\varnothing)$ and $v_{\theta}(x|y)$ are the unconditional and conditional predictions, and $w$ is guidance scale. In diffusion models, inference with high guidance scale is stable because latent variables are typically normalized, ensuring that conditional and unconditional predictions maintain a consistent scale. However, in token-level autoregressive models, global normalization of the entire latent tensor does not enforce per-token statistical consistency. Consequently, small discrepancies between conditional and unconditional predictions are magnified by a large guidance scale, leading to a significant drift in the statistics of generated tokens over the sequence.
@@ A Regularized Latent Space is Critical for Generation
Figure 7: Latent distributions in 16 channels for three VAE variants: Flux.1-dev, NextStep-1 w/o noise, and NextStep-1. Blue bars show empirical histograms; red lines indicate the standard normal distribution. NextStep-1 VAE aligns best with the normal distribution, reflecting dispersed latent distribution.
@@
We attribute this phenomenon to noise regularization cultivating a well-conditioned latent space. This process enhances two key properties: the tokenizer decoder’s robustness to latent perturbations (Figure˜6) and a more dispersed latent distribution (Figure˜7), a property prior work has also found beneficial for generation (Yang et al., 2025; Yao et al., 2025; Sun et al., 2024c). While it remains unclear whether robustness or dispersion plays a critical role, these results underscore the practical benefits of noise-based regularization and highlight promising directions for future analysis.
@@ 6.3 Limitations and Challenges
While NextStep-1 successfully demonstrates that autoregressive models can operate on high-dimensional continuous latent spaces, achieving generation quality comparable to diffusion models, this approach also introduces unique stability challenges. We observed the emergence of several distinct generative artifacts when transitioning from a VAE with a lower-dimensional latent space (e.g., spatial downsample factor is 8 and number of latents channel is 4) to one with a higher-dimensional space (e.g., spatial downsample factor is 8 and number of latents channel is 16). While the former configuration produced stable outputs, the latter occasionally exhibited failure modes, as illustrated in Figure˜8.
@@
Our framework faces two primary challenges in scaling to high-resolution image generation, particularly when compared to diffusion models, which benefit from well-established techniques (Esser et al., 2024; Chen et al., 2024) in this domain. First, the strictly sequential nature of autoregressive generation requires substantially more training steps to converge at higher resolutions. In contrast, diffusion models refine the entire image in parallel at each iteration, enabling more direct exploitation of 2D spatial inductive biases. Second, techniques recently developed for high-resolution diffusion models, such as timestep shift, are difficult to adapt to our setting. This limitation arises because the Flow Matching Head acts primarily as a lightweight sampler, while the transformer backbone performs the core generative modeling; thus, modifications to the sampling process have only marginal impact on the final output. Designing high-resolution generation strategies specifically for patch-wise autoregressive models remains an important direction for future research.

### v2

@@ 1 Introduction
However, most existing autoregressive approaches for text-to-image generation (Sun et al., 2024b; Wang et al., 2024b; Sun et al., 2024a; Chen et al., 2025b; Tong et al., 2024; Dong et al., 2024) either rely on heavy diffusion models or adopt vector quantization (VQ) (Zheng et al., 2022; Eslami et al., 2021; Yu et al., 2023) to tokenize images into discrete visual tokens, which encounter limitations including exposure bias (Han et al., 2025) and suboptimal image tokenization (Li et al., 2024c). While recent efforts with continuous latent representations (Li et al., 2024c; Fan et al., 2024; Sun et al., 2024c; Tschannen et al., 2025, 2024) have shown promise, a significant performance gap persists between autoregressive models and state-of-the-art diffusion methods (Podell et al., 2024; Esser et al., 2024; Labs, 2024), particularly in image quality and consistency.
@@ 2.2 Model Architecture
The latents from the image tokenizer are pixel-shuffled into a more compact sequence. This is achieved by applying a space-to-depth transformation with a $\times2$ kernel, which flattens $\times2$ spatial latents into the channel dimension. For example, this converts the latents of a $\times256$ image into a $\times16$ grid of 64-channel tokens. This grid is then flattened into a 1D sequence of 256 tokens to serve as input for the following Causal Transformer.
@@ 3.2 Image-Text Pair Data
    Data Sourcing: We collected a large-scale dataset from diverse sources, including web data, multi-task VQA data, and text-rich documents.
@@ 3.3 Instruction-Guided Image-to-Image Data
For visual perception and controllable image generation tasks, we synthesized 1M samples by applying the annotator of ControlNet (Zhang et al., 2023b) to a part of our high-quality image-text pair data. For image restoration and general image editing, we collected 3.5M samples, comprising data from GPT-Image-Edit (Wang et al., 2025d), Step1X-Edit (Liu et al., 2025), and a proprietary in-house dataset. Following Step1X-Edit (Liu et al., 2025), all editing data were subjected to a rigorous VLM-based filtering pipeline that assessed both image-pair quality, rationality, consistency, and instruction alignment, resulting in about 1M high-quality instruction-guided image-to-image data for training.
@@ 4.1 Training Image Tokenizer
Our image tokenizer is initialized from the Flux.1-dev VAE (Labs, 2024), selected for its strong reconstruction performance. We fine-tune this model on the image-text dataset detailed in Section˜3.2 to adapt it to our specific data distribution. For optimization, we employ the AdamW optimizer (Loshchilov and Hutter, 2019) with $(\beta_{1}=0.9,\beta_{2}=0.95,\varepsilon=1\times 10^{-8})$ for its convergence stability. The model is trained for 50K steps with a total batch size of 512, using a constant learning rate of $1\times 10^{-5}$ preceded by a linear warm-up of 1000 steps.
@@ 4.3 Post-Training
PixArt-Alpha (Chen et al., 2024b) 0.48 - - 71.11
@@ 5.1 Performance of Text-to-Image Generation
PixArt-Alpha (Chen et al., 2024b) 0.45 0.50 0.48 0.49 0.56 0.34 0.47 0.63
@@
GPT-Image-Edit (Wang et al., 2025d) - - 7.24 - - - 3.80
@@ 6.1 What Governs Image Generation: the AR Transformer or the FM Head?
A key architectural distinction of our framework is its direct, autoregressive modeling of continuous image tokens using a flow matching objective. Prevailing autoregressive models for image generation  (Sun et al., 2023, 2024b; Dong et al., 2024; Zhou et al., 2025; Chen et al., 2025a) typically rely on heavy diffusion models for an entire image: an autoregressive model first produces a semantic embedding, which is then used to condition a diffusion model that generates an entire image in a single denoising process. In contrast, our model autoregressively generates the image patch-by-patch (Kou et al., 2024; Chen et al., 2024a; Fan et al., 2025), modeling the distribution of each patch with a significantly more lightweight flow matching model. We argue that this establishes our framework under the pure autoregressive paradigm with next-token prediction (NTP) modeling, rather than a diffusion model merely orchestrated by a Transformer.
@@ 6.2 Tokenizer is the Key to Image Generation
where $v_{\theta}(x|\varnothing)$ and $v_{\theta}(x|y)$ are the unconditional and conditional predictions, and $w$ is guidance scale. In diffusion models, inference with a high guidance scale is stable because latent variables are typically normalized, ensuring that conditional and unconditional predictions maintain a consistent scale. However, in token-level autoregressive models, global normalization of the entire latent tensor does not enforce per-token statistical consistency. Consequently, small discrepancies between conditional and unconditional predictions are magnified by a large guidance scale, leading to a significant drift in the statistics of generated tokens over the sequence.
@@ A Regularized Latent Space is Critical for Generation
Figure 7: Latent distributions in 16 channels for three VAE variants: Flux.1-dev, NextStep-1 w/o noise, and NextStep-1. Blue bars show empirical histograms; red lines indicate the standard normal distribution. NextStep-1 VAE aligns best with the normal distribution, reflecting a dispersed latent distribution.
@@
We attribute this phenomenon to noise regularization, cultivating a well-conditioned latent space. This process enhances two key properties: the tokenizer decoder’s robustness to latent perturbations (Figure˜6) and a more dispersed latent distribution (Figure˜7), a property prior work has also found beneficial for generation (Yang et al., 2025; Yao et al., 2025; Sun et al., 2024c). While it remains unclear whether robustness or dispersion plays a critical role, these results underscore the practical benefits of noise-based regularization and highlight promising directions for future analysis.
@@ 6.3 Limitations and Challenges
While NextStep-1 successfully demonstrates that autoregressive models can operate on high-dimensional continuous latent spaces, achieving generation quality comparable to diffusion models, this approach also introduces unique stability challenges. We observed the emergence of several distinct generative artifacts when transitioning from a VAE with a lower-dimensional latent space (e.g., spatial downsample factor is 8 and number of latent channels is 4) to one with a higher-dimensional space (e.g., spatial downsample factor is 8 and number of latent channels is 16). While the former configuration produced stable outputs, the latter occasionally exhibited failure modes, as illustrated in Figure˜8.
@@
Our framework faces two primary challenges in scaling to high-resolution image generation, particularly when compared to diffusion models, which benefit from well-established techniques (Esser et al., 2024; Chen et al., 2024b) in this domain. First, the strictly sequential nature of autoregressive generation requires substantially more training steps to converge at higher resolutions. In contrast, diffusion models refine the entire image in parallel at each iteration, enabling more direct exploitation of 2D spatial inductive biases. Second, techniques recently developed for high-resolution diffusion models, such as timestep shift, are difficult to adapt to our setting. This limitation arises because the Flow Matching Head acts primarily as a lightweight sampler, while the transformer backbone performs the core generative modeling; thus, modifications to the sampling process have only a marginal impact on the final output. Designing high-resolution generation strategies specifically for patch-wise autoregressive models remains an important direction for future research.